• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Gao

    Log entry time 00:16:24 on June11,2009

    Entry number 280504

    keyword=Swing shift summary

    June 10 Swing Shift
    Started with run 3448 and 22319, 7M ended 16:10 pm
    16:10 Target NMR measurement 60%
    Started 3449, 22320
    16:45 stopped run 3449 and 22320
    Yi turned off all lasers with Vince's approval
    because the connectors between fibers got rather hot
    Started a new Compton run 17303
    16:51 started run 3450 and 22321
    Yi noticed that one of the fibers was damaged and is going to replace it.
    17:25 stopped run 3450 and 22321
    started run 3451 and 22322
    17:55 called MCC to insert beam half-wave plate
    Started run 3452, 22323 and Happex 31508
    18:10 Yi started to turn the lasers back on
    18:25 Lasers tripped off
    18:35 started run 3453 and 22324 (Yi is working on the lasers again)
    18:45 Target Pol. measurement 51.2%
    19:05 started runs 3454 and 22325
    19:35 started runs 3455 and 22326
    20:05 started runs 3456 and 22327
    20:40 started runs 3457 and 22328 New Compton run 17304
    21:10 Started runs 3458 and 22329
    21:50 Started runs 3459 and 22330
    22:20 started runs 3460 and 22331 New Happex 31509
    22:51 started runs 3461 and 22332 New Compton 17306
    23:20 NMR measurement of target polarization 60.4%
    23:25 started runs 3462 and 22333
    23:50 started runs 3463 and 22333